OpenStudy (sleepyjess):

Find an explicit rule for the nth term of a geometric sequence where the second and fifth terms are -36 and 2304, respectively.

OpenStudy (sleepyhead314):

I just tried teaching this to myself on the spot, so I might be wrong. let X be the equation you want to find. X = a r^(n-1) is the geometric sequence equation where a is the first term r is the rate and n is the output term @sleepyjess

OpenStudy (sleepyhead314):

so with this, we can get the two equations a r^(2-1) = -36 and a r^(5-1) = 2304 do you see where I got that? :)

OpenStudy (sleepyhead314):

we can simplifyish those two into a r^1 = -36 and a r^4 = 2304 just simplifying the exponent part ;) now this is like the substitution solving method idea take the first equation, solve for a we get a = -36 / r right? then let's substitute it into the second equation :) [(-36)/r]*r^4 = 2304 \[\frac{ -36 }{ r }*r^4 = 2304\] solve for r

OpenStudy (sleepyhead314):

once you get r plug it back into a = -36 / r = ? then plug all that back into X = a r^(n-1) so you would end up with X = (#)*(#)^(n-1) <-- as in keep the n variable in your final answer

OpenStudy (sleepyhead314):

source: example 2 http://hotmath.com/hotmath_help/topics/nth-term-of-a-geometric-sequence.html

OpenStudy (sleepyjess):

Sorry, there was an armadillo in our garage....

OpenStudy (sleepyhead314):

take a look at what I've written and tell me if you're still confused ^_^

OpenStudy (sleepyjess):

kinda

OpenStudy (sleepyhead314):

lol what part? I can try to explain again

OpenStudy (sleepyhead314):

@iambatman I feel like I might be doing this wrong? or no?

OpenStudy (sleepyjess):

I don't really get any of it :/

OpenStudy (sleepyjess):

@e.mccormick

OpenStudy (sleepyhead314):

read through what I typed again carefully and try to get it I'm out for the night and will hand you the answer tomorrow morning if you still don't get it